OpenStudy (anonymous):

how to find the missing coordinate given a slope of 1/2 and the points (3, -12) and (x, -18)

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) 1/2=(-18+12)/(x-3) 1/2=(-6)/(x-3) x-3=-6*2 x-3=-12 x=-12+3 x=-9